Rebel Sole is a rapidly expanding shoe company. The following is the demand estimate for its popular shoes. The estimate was done using 40 observations.

Q = 10 – 10 P + 4 A + 0.42 I + 0.25 Py

       (3)   (1.8)    (0.7)   (0.1)     (0.1)

F = 93, s = 6, R2 = 93%

Q is quantity sold (in thousands), P is shoe price, A is advertising expenditure (in thousands), the numbers in parentheses are standard errors, I is disposable income per capita (thousands of dollars), and Py is the price of related goods.

  1. Evaluate the model based on F, R2.
  2. Test the significance of Py.
  3. If P = $5, A = $30,000, I = 50,000, and Py = $6, calculate advertising elasticity.
  4. Given the information in c. above, calculate the 95% confidence interval for Q.
Answer #1

(a)

F = 93

All independent variables have a jointly significant effect on Q.

R2 = 93%

Represents that price, advertising, disposable income and price of related goods cause 93% variability in Q.

(b)

Z = 0.25 / 0.1 = 2.5

Z is 1.96 for 5% of level significance, which is less than 2.5.

Explanation: Py is significant.

(c)

Q = 10 -10P + 4A + 0.42I + 0.25Py

Q = 10 – 10(5) + 4(30) + 0.42(50) + 0.25(6)= 10- 50 + 120 + 21 +1.5

= 102.5

Advertising Elasticity = 4 x 30 / 102.5 = 1.17

(d)

Q.= 0.1 ± 1.96 (3) = 0.1 ± 5.88

Lower Limit= - 5.78

Upper Limit = 5.98